OpenStudy (anonymous):

A boat can maintain a constant speed of 16mph relative to the water. The boat makes a trip upstream in 20min and the return trip takes 15min. What is the speed of the current?

OpenStudy (anonymous):

Solve the following for c, the speed of the current:\[(16-c)\frac{20}{60}=(16+c)\frac{15}{60} \]\[c\text{ = }\frac{16}{7}=2.29\text{ mph}\]